100 kg/min of steam at 34 bars, 340°C (h=3082.0kJ/kg) expands through a turbine isentropically until the pressure becomes 0.25 bars (h= 2540.37kJ/kg). Find the work in hp.
1
Expert's answer
2020-10-22T18:28:17-0400

Work done in isentropically process is given by, "W = m(h_1 - h_2)"

So, Putting values,


"W = 100\times (3082.0 - 2540.37) = 54163 KJ"